OpenStudy (anonymous):

Carol has 3 children. The youngest child is 4 years younger than the middle child. The oldest child is 6 years older than the middle child. The youngest child is 8. The sum of the their age is 12 years less than Carol's age. How old is Carol?

OpenStudy (anonymous):

do they give u choses

OpenStudy (anonymous):

so the proble is going to have some steps

OpenStudy (anonymous):

oldest child= m+6 middle child = ? youngest child= m-4=8 This is what we start off knowing. So then we find "m" to solve. oldest = 18 middle = 12 youngest = 8 All together is 38 So 38-12= Carol's Age Carol = 26 (pretty young to have an 18 year old kid though...)

OpenStudy (anonymous):

a) c = y-(y-4)-(y-4-6)-12 b) c= y+(y+4)+(y+4+6)+12 c) c+12=y+(y+4)+(y+4+6 d) c = y +(y -4) +(y-4-6)+12 these are my answer choices

OpenStudy (anonymous):

* Sorry I meant 38+12 Carol= 50 years old!

OpenStudy (anonymous):

they are asking for a formula im sorry i forgot to mention